OpenStudy (anonymous):

Solve the equation, where 0° ≤ x < 360°. Round approximate solutions to the nearest tenth of a degree. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.) 3 sin x − 2 cos x = 1 ..please solve.help me

OpenStudy (xapproachesinfinity):

hmm this is gonna require some attention first i would want you to think of asin(x+b)=asinxsinb+acosxcosb

OpenStudy (xapproachesinfinity):

I want to achieve 3sinx-2cosx okay

OpenStudy (xapproachesinfinity):

if i set my first expression equal to that i would get asinb=3 and acosb=-2

OpenStudy (xapproachesinfinity):

i need here find a first

OpenStudy (xapproachesinfinity):

now we need to recall the equation of a circle x^2+y^2=r^2 my r here is a x=-2 and y=3 so \(r=\sqrt{(-2)^2+3^2}=\sqrt{13}\)

OpenStudy (xapproachesinfinity):

i mean \(a=\sqrt{13}\) now we need the angle b

OpenStudy (xapproachesinfinity):

so any idea how to get b?

OpenStudy (xapproachesinfinity):

@blowfish

OpenStudy (xapproachesinfinity):

i need tanb here i have tanb=sinb/cosb=-3/2 you can use arctan(-3/2)=? to get the angle and then go back to the original equal that i wrote asin(x+b)=1 substitute a and b by the values you found and then solve

OpenStudy (xapproachesinfinity):

I think you got a pretty good idea here

OpenStudy (anonymous):

I haven't got the right answer

OpenStudy (xapproachesinfinity):

what did you do?

OpenStudy (xapproachesinfinity):

Did you follow what i did

OpenStudy (anonymous):

I squared both sides and substitute it in the quadratic formula

OpenStudy (xapproachesinfinity):

hmm squared both sides i don't think that would work

OpenStudy (xapproachesinfinity):

show me what you did

OpenStudy (anonymous):

what am I going to do, all I input in my quiz online were wrong.. will you solve it for me..what are the angles

OpenStudy (anonymous):

(3sinx)^2=(2cosx+1)

OpenStudy (anonymous):

(2cosx+1)^2

OpenStudy (xapproachesinfinity):

hmmm so what did you get after that

OpenStudy (anonymous):

-2+or-2square root of 11 over 5

OpenStudy (xapproachesinfinity):

hmmm i don't think that gonna work pal? you will be circulating for ever around to find something hehe

OpenStudy (anonymous):

so can you find the angles?

OpenStudy (xapproachesinfinity):

i give you the method to get them

OpenStudy (xapproachesinfinity):

read my comments carefully do what i suggested

OpenStudy (anonymous):

I don't know how to manipulate it.. sorry

OpenStudy (anonymous):

can someone help me pls.I tried it but can't find answers

OpenStudy (xapproachesinfinity):

Did you read what idid above from a to z i give you everything you need to get this

OpenStudy (xapproachesinfinity):

Reread it again with attention

OpenStudy (anonymous):

YEah I read it but really can't find the answer everytime I input my answer online

OpenStudy (anonymous):

well, anyways, thank you so much for your effort but my assignment had closed already...

OpenStudy (xapproachesinfinity):

i don't think you read it i said find b in asin(x+b)=1 tanb=-3/2 ====> artan(-3/2)=? use the calc to find b

OpenStudy (xapproachesinfinity):

a=root(13) i showed how we got it

OpenStudy (xapproachesinfinity):

hehe there was a time limit i see Quiz then! My goal is to assist you not giving you direct answers but at any rate me too i need a calculator to find b because it is a decimal angle you can't do this in mind hehe